OpenStudy (anonymous):

Differentiate the following function. f(t)=8(sqrt t)-(8)/(sqrt t)

OpenStudy (nowhereman):

These are just the basic differentiation rules. Do you know the derivative of the square root?

OpenStudy (anonymous):

Yea so far i have 8(1/2 t^-1/2)-8(/1-2 t ^-1/2) now im stuck on that location

OpenStudy (nowhereman):

This is really hard to read. When you have \(1/\sqrt t\) you need to deal with it like \(t^{-\frac{1}{2}}\) so you'll get \(-\frac{1}{2} t^{-\frac{3}{2}}\)

OpenStudy (anonymous):

im not understanding what you are writing theres too many //[[{{

OpenStudy (nowhereman):

Then you should turn on javascript, so that it can get rendered as latex.

OpenStudy (anonymous):

Yeah, make sure you have javascript on and then it'll look alot better :P

OpenStudy (anonymous):

i am but its taking to long

OpenStudy (anonymous):

\[4/\sqrt{t} + 4/t^{3/2}\]

OpenStudy (anonymous):

ok thanks SmRoberts..im still trying to figure out where i got wrong and my java is not working at the moment

OpenStudy (anonymous):

happy to help

OpenStudy (anonymous):

df(t)/dt=d/dt[8(sqrt t)-(8)/(sqrt t)] =8(1/2 t^-1/2)+8(1/2 t ^-3/2) =(4 t^-1/2)+(4 t ^-3/2) answer

OpenStudy (anonymous):

did u get it rmalik...